Mfmod pcre

version 1.1, 26 June 2023

Sergey Poznyakoff.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2022–2023 Sergey Poznyakoff Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover, and no Back-Cover texts. A copy of the license is included in the section entitled “GNU Free Documentation License.”

Table of Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Basic Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3 Matching Entire Messages . . . . . . . . . . . . . . . . . . . . . . 2 4 Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

4.1 Compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2 Matching Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.3 Working with Capturing Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.4 Message Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5 Downloads and Other Links . . . . . . . . . . . . . . . . . . . . 6 6 Bug Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Appendix A GNU Free Documentation License . . 6 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

1 Overview

This package extends mailfromd, a powerful mail filtering utility designed for use with mail transport agents supporting the milter protocol (see Mailfromd Manual).

Mail filtering language (MFL) which is used to program mailfromd, uses POSIX regular expressions (either basic or extended, at the user’s option). Mfmod_pcre extends it by providing support for Perl-compatible regular expressions.

It is written as a loadable module (see Section “mfmod” in Mailfromd Manual) and requires mailfromd version 8.14.91 or later and pcre2 10.35 or later.

This manual describes mfmod_pcre version 1.1 and discusses its use. It supposes that the reader is sufficiently acquainted with mailfromd and Perl-compatible regular expressions.

For a complete description of mailfromd, refer to Mailfromd Manual.

For a description of Perl-compatible regular expressions, see https://pcre.org/

original/doc/html/pcresyntax.html.

2 Basic Usage

To load mfmod_pcre into your MFL program, first make sure that mailfromd search path (see Section “include search path” in Mailfromd Manual) contains directory where the pcre.mf interface file is installed and search path for dynamically loaded modules (see Section “mfmod-path” in Mailfromd Manual) contains directory where the loaded library mfmod_pcre.so resides. Once these prerequisites are met, add the require statement:

     require 'pcre'

Now you can use Perl-compatible regexps in your code.

Names of all functions exported by the module begin with the pcre_prefix. For example, pcre_string_matches checks whether its second argument matches a regular expression supplied as its first argument. The function returns true (1) if the string matches the expression, and false (0) otherwise. E.g.:

     if pcre_string_matches('.+?@.+', address)
       # do something with address
     fi

This function suits for casual use, but you should be aware of its drawbacks. First, it compiles regular expression on each call, which hurts performance. Secondly, while it allows to use capturing groups in regex, it provides no way to retrieve matched substrings.

The function pcre_string_match is free from the second limitation. If the string matches, this function returns a match descriptor, a positive integer value describing the match. This value can then be used with other functions to retrieve information about the match. For example, pcre_numbered_substring retrieves the captured string:

     set mh pcre_string_match('(\d+)\s+(.+?)\s+(\S+@\S+)', input_line)
     if mh
       # Regular expression matches
       echo "Ordinal number: " . pcre_numbered_substring(mh, 1)
       echo "Personal part: " . pcre_numbered_substring(mh, 2)
       echo "Email: " . pcre_numbered_substring(mh, 3)
       pcre_match_free(mh)
     else
       # Expression doesn't match
     fi

Notice the use of pcre_match_free: when no longer needed, match descriptor should be disposed of using this function.

A similar function pcre_named_substring is provided to retrieve named captured groups.

Both numbered and named groups can be expanded within a string using the pcre_ expand function. Numbered groups are referred to as $n and named ones as $+{name}. For example:

     set mh pcre_string_match('(?<ord>\d+)\s+(?<personal>.+?)\s+(?<email>\S+@\S+)', input_line)
     if mh
       # Regular expression matches
       set printable_email pcre_expand(mh, '"$+{personal}" <$+{email}>')
       pcre_match_free(mh)
     fi

However, pcre_string_match still suffers from the first limitation: it compiles the regular expression on each call. To overcome this limitation, use precompiled regular expressions.

The function pcre_compile compiles the regexp given as its argument and returns regexp descriptor, a numeric value that is used to refer to that expression in other pcre_calls:

     number rx pcre_compile('(\d+)\s+(.+?)\s+(\S+@\S+)')

Thus, a regexp can be compiled once (e.g. in the begin section) and then used as many times as needed. A set of functions is provided to use precompiled regexps. For example, pcre_matches and pcre_match are similar to the functions discussed above, except that they take regexp descriptor as their first argument:

     number rx pcre_compile('(\d+)\s+(.+?)\s+(\S+@\S+)')
     set mh pcre_match(rx, input_line)
     if mh
       # Regular expression matches
       set printable_email pcre_expand(mh, '"$2" <$3>')
       pcre_match_free(mh)
     fi

3 Matching Entire Messages

Two functions are provided to search for regular expressions within emails. They are declared as follows:

     func pcre_match_message(number rx, number msg; string charset)
       returns string
     func pcre_string_match_message(string regex, number msg; string charset)
       returns string

The message to search in is given by message descriptor, i.e. an integer number identifying it. This descriptor is returned by one of the following calls:

mailbox_get_message

See Section “Mailbox functions” in Mailfromd Manual.

current_message

See Section “current message” in Mailfromd Manual.

message_from_stream

See Section “Message functions” in Mailfromd Manual.

Message searching fuctions can be used to search in plain as well as in multipart (MIME) messages. When searching in a MIME message, it is descended recursively and each its part is searched individually.

On success (a match is found), these functions return a string describing where the match was found. For example, ‘1’ means that the plaintext message body matched, ‘1.2’ means second part of a multipart message, ‘1.2.3’ means third subpart of the second part of a multipart message, etc.

On error, an empty string is returned.

Only textual parts of a MIME message are considered. A textual part, for the matching purposes, is a part whose Content-Type header matches one the globbing patterns in the text mime type list. By default this list contains a single value: ‘text/*’.

More patterns can be added to text mime type list using the pcre_text_mime_add function. It takes variable number of arguments, e.g.:

     pcre_text_time_add("application/x-json", "x-text/*")

Similarly, the function pcre_text_mime_del allows you to delete elements from this list. Note, that it treats its arguments literally, not as globbing patterns. For example, suppose the following call is done when the text mime type list contains two elements: ‘text/*’,

and ‘*/x-sh’:
     pcre_text_time_del('text/*')

This will remove only one item: ‘text/*’, but won’t touch ‘*/x-sh’ (which it would do, if it treated its argument as a globing pattern).

Finally, the function pcre_text_mime_init removes all entries from the list.

4 Functions

This chapter describes function calls exported by the module. The functions are grouped by their purpose.

4.1 Compile

[Function] number pcre_compile (string pattern)

Compiles regular expression pattern and returns regexp descriptor, a numeric value that will be used to refer to that regular expression when calling other pcre_functions. If the regular expression is erroneous and cannot be compiled, the function throws e_regcomp exception (see Section “Built-in Exceptions” in Mailfromd Manual).

4.2 Matching Strings

The functions below return match descriptor, i.e. a positive number that can be used to retrieve further information about the match (see Section 4.3 [Capturing groups], page 4). On failure (string doesn’t match the oregular expression), they return 0.

If the supplied match descriptor is not valid, these functions will raise the e_range exception.

[Function] number pcre_match (number rx, string subject)

First argument is regex descriptor returned by previous call to pcre_compile. The function matches subject against the regular expression referred to by rx. Returns match descriptor on success and 0 on failure.

[Function] number pcre_string_match (string rs, string subject)

Compiles rs into a regular expression and matches string subject against it. Returns match descriptor on success and 0 on failure.

When no longer used, the match descriptor should be freed:

[Function] void pcre_match_free (number mh)

Frees the resources allocated for the match descriptor mh and returns it to the pool for eventual re-use.

The two functions below return a boolean value: true (1) if the string matches and false (0) if it doesn’t.

[Function] number pcre_matches (number rx, string subject)

First argument is regex descriptor returned by a previous call to pcre_compile. Returns true if string matches that regular expression and false otherwise.

[Function] number pcre_string_matches (string rs, string subject)

Compiles rs into a regular expression and matches string subject against it. Returns true if the match is found and false otherwise.

4.3 Working with Capturing Groups

The following functions are used to retrieve actual content of a captured group after a successive match. The match is identified by match descriptor returned from a prior call to

prce_match or pcre_string_match.

These functions will raise the e_range exception, if the match descriptor is invalid, and the e_inval exception, if the substring number or name refer to a non-existing capturing group.

[Function] string pcre_numbered_substring (number mh, number n)

Returns text of the nth captured group of the match mh.

[Function] string pcre_named_substring (number mh, string name)

Returns text of the named captured group name in the match mh.

[Function] string pcre_expand (number mh, string input)

Given match descriptor mh, replace in input all references to capturing groups with their actual text and return the resulting string. Numbered capturing groups are referred to as ‘$n’, where n is the group number (0-based, 0 being the portion of the text matching entire regular expression). If n is greater than 9, it should be enclosed in curly braces (e.g. ‘${10}’). Named capturing groups are referred to as ‘$+{name}’, where name is the group name.

4.4 Message Matching

The functions in this section match the body of an email message against a regular expression. If the message is a multipart one (MIME), then its parts are descended recursively and each part that is recognized as a textual one is matched against the regexp.

The message is identified by its descriptor, a numeric value returned from a previous call

to mailbox_get_message, message_from_stream, current_message, or similar function.

On success, these functions return a string identifying the part of the message that matched. For example, ‘1’ means the body of a plaintext email message. ‘1.2’ means second part (part numbers are 1-based) of a multipart message.

If the match isn’t found, empty string is returned.

[Function] string pcre_match_message (number rx, number msg; string

   charset)

Checks whether the body of the message msg (or any of its parts, if it is a MIME message) matches regular expression referred to by descriptor rx. If optional charset is supplied, each message part is re-coded to that charset before matching.

[Function] string pcre_string_match_message (string rs, number msg; string
   charset)

Compiles rs and checks whether the body of the message msg (or any of its parts, if it is a MIME message) matches it. If optional charset is supplied, each message part is re-coded to that charset before matching.

These functions use the Content-Type header to decide whether to apply the regular expression to the given message part, or just to skip it. An internal structure called text mime type list contains globbing patterns of the eligible content type values. The actual value of the Content-Type header is matched against each list element in turn, until either a match is found (in which case the part will be processed) or the end of list is reached (in which case the part will be skipped).

By default, the text mime type list is initialized with a single value: ‘text/*’. To maintain the list, the following functions are provided:

[Function] void pcre_text_mime_init ()

Clears the list.

[Function] void pcre_text_mime_add (...)

Arguments are strings in form ‘part/subpart’, where both part and subpart can be shell-style globbing patterns. The supplied strings are appended to the list.

[Function] void pcre_text_mime_del (...)

Deletes from the list those patterns that match literally its arguments.

5 Downloads and Other Links

The program can be downloaded from https://download.gnu.org.ua/release/
mfmod_pcre.
The source repository is available at https://git.gnu.org.ua/mfmod_pcre.git/.
The package development page is at https://puszcza.gnu.org.ua/projects/
mfmod_pcre.

6 Bug Reports

If you think you found a bug in mfmod_pcre or in its documentation, please send a mail to gray@gnu.org (Sergey Poznyakoff) or use the bug tracker at https://puszcza.gnu.org.

ua/bugs/?group=mfmod_pcre (requires authorization).

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 https://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers.

The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work.

In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS

You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects.

You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS

A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document.

If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION

Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail.

If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See https://www.gnu.org/licenses/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization.

“Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

Index 14

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Index

B

boolean matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

C

capturing group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

E

e_inval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 e_range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 e_regcomp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

M

mailfromd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 match descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 4 matching boolean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 message descriptor. . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 5 MFL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 mfmod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 module, using . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

P

pcre . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 pcre_compile, example . . . . . . . . . . . . . . . . . . . . . . . . . 2 pcre_expand, example . . . . . . . . . . . . . . . . . . . . . . . . . . 2 pcre_match_free, introduced . . . . . . . . . . . . . . . . . . . 2

                                                pcre_numbered_substring, example. . . . . . . . . . . . . 1

pcre_string_matches, example . . . . . . . . . . . . . . . . . 1 pcre2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

R

regexp descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 4

S

search and replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

T

text mime type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 text mime type list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 textual part defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3